Hello, I am enrolled in College Algebra, and I am faced with a problem involving Population, Specifically:
In 200, the population of Greece was 10,600,000, with projections of a population decrease of 28,000 people per year. In the same year, the population of Belgium was 10,200,000, with projections of a population decrease of 12,000 people per year.
This problem requires two solutions, and has to be solved algebraically, showing the equation used and work used to achieve the answers.
a) When will the two countries have the same population?

Equation used:___________ & What year?
b) What will be the population at that time? & Population size:

Answer by ankor@dixie-net.com(22740)

In 200, the population of Greece was 10,600,000, with projections of a population decrease of 28,000 people per year.
In the same year, the population of Belgium was 10,200,000, with projections of a population decrease of 12,000 people per year.
This problem requires two solutions, and has to be solved algebraically, showing the equation used and work used to achieve the answers.
:
let t = the number of year for this to be true, from the year 2000 (not 200)
Write an equation for each country
Greece: f(t) = -28000t + 10600000
Belgium: f(t) = -12000t + 10200000
:
a) When will the two countries have the same population?
when these two function are equal
hence the equation:
-12000t + 10200000 = -28000t + 10600000
some basic algebra here
-12000t + 28000t = 10600000 - 10200000
+16000t = 400000
t = 400000/16000
t = 25 years, we can say they should be equal in 2025
:
:
b) What will be the population at that time? & Population size
Greece: -28000(25) + 10600000 = 9,900,000
Belgium: -12000(25) + 10200000 = 9,900,000